OMERO is a piece of software which is part of the Open Microcopy Environment. It is specially designed for the management of biological microscopy data, in other words; images. The facility provides this software as part of their services.

This system should not be used for highly sensitive data, for example patient data, etc. All other scientific data is safe with us.

Important notices

imaris-2 New Imaris version 9.6.0. Updated Baramundi jobs for Imaris and ImarisViewer. Put the new installers on

leica-tcs-sp5-med Z motor of the objective turret does not work properly. It cannot assume the usual lowest position and makes strange noises.

leica-tcs-sp5-bio Training Celine Talon (Dengjel Group)

leica-dm6b-navigator Training:, (Walch group) no

leica-tcs-sp5-med Fixed screen scaling issue for users that login the first time. no

leica-tcs-sp5-bio Fix the custom scaling issue for users that login the first time. no

leica-tcs-sp5-bio Installed OMERO.insight 4.10 and Java 8. (running the 32 bit version) no

leica-tcs-spe-ii-bio Fix screen scaling issue for fist logins of users: no

leica-tcs-spe-ii-bio Installed OMERO.insight 4.10 no

leica-dm6b-navigator training Blanchod group (3 students) no

leica-tcs-spe-ii-bio training: blanchod group (2 students) no

visitron_csu-w1 Error message appears when using FRAP Modul to set a region. Overflow Error: Array dimensions exceeded supported range. line 5

leica-tcs-sp5-med Replaced mercury lamp 499-28 (Fluo for binoculars) no

huygens-2 Finally found why my personal account did not work anymore: the inc/Validator.php has a pwd check that invalidates strings containing white spaces. The condition was commented out and I wrote an issue on the HRM dev issue tracker. no

huygens-2 There is currently an issue with the SSL handshake of the late OMERO 5.4 minor versions, the workaround is to use: omero import --skip upgarde. The was adapted accordingly on the deployed-2 branch no

huygens-2 Updated HRM to 3.7.0 and Huygens Core to 19.10. no

leica-tcs-sp5-med DAPI does not work with binocculors. Fluo bulb is not working. no

leica-tcs-sp5-bio Setup Backup. Setup the script to launch bioimage website at startup. Put HRM shortcut on the Desktop for all user. no

leica-tcs-sp5-bio Connect the new computer to the UNIFR domain. no

leica-tcs-sp5-med Setup the Backup, HRM shortcut on the desktop, install script to launch the bioimage website. no

leica-tcs-sp5-med Connected the computer to the UNIFR domain. no

leica-tcs-spe-ii-bio Setup backup, add HRM shortcuts to the Desktop, install startup script to launch the bioimage website. no

leica-tcs-spe-ii-bio Added the computer the to UNIFR domain no

leica-tcs-sp5-bio Beamsplitter was replaced by Leonardo Silva and is functional.

leica-tcs-sp5-med New PC workstation and upgrade to Windows10. Needs to added to the network

leica-tcs-spe-ii-bio New PC workstation and upgrade to Windows10. Needs to be added to the network. no

leica-tcs-sp5-bio System is unstable, software freezes. Workstation needs to be added to the network to arrange for a Leica remote session

ge-deltavision-elite-med Maintenance and TIRF adjustments: - 514 Laser does not have enough power anymore (bellow 1 mW), all the other lasers are strong enough. - TIRF arm was aligned (prism had to be turned) - The RDC connection to the mic. controller did not work initially. After doing a new 'shortcut' on the desktop it worked again (probably is a script we should look at) - 60x TIRF objective installed and tested. no

omero-2 PathViewer upgrade to 3.0.4. The interface changed considerably. Navigation is faster. The upgrade broke the links used in the slide-descriptions. no

omero-2 Debian upgrade (minor) on bioimage-db and virutual-microscopy no

omero-2 OMERO minor upgrade to 5.4.10 (bioimage-db and virtual-microscopy, no upgrade necessairy for HRM OMERO integration) no

leica-dm6b-navigator training: no

ge-deltavision-elite-med Training Ola El Atab (Schneiter Group) no

leica-tcs-sp5-bio Training Marta Sprecher (Sprecher Group) and Isabel van der Decken (Lauber Group) no

leica-tcs-sp5-bio The motor of the beam splitter to change the dichroic mirrors seems to respond only very slowly. As a consequence fast multi-channel acquisitions by using 405nm laser line in sequential mode with other lines are not possible. Hence we recommend Users to use the TCS Leica SP5 (Med) system for fast multi-channel acquisitions. To work around this problem on the TCS Leica SP5 (Bio) avoid switching the dichroic mirror when using sequential acquisition mode. We recommend to use the triple dichroic mirror (TD) for 488/543/633nm laser lines and switch to 405nm line at last only, preferentially using stack mode (not line or frame mode). yes

hamamatsu-nanozoomer-med Fluo-calibration done. The stripes and shading is gone. However, a test scan yielded constantly saturated DAPI signal. The light source had to be tuned down to 40% light intensity so that the image were not saturated.

hamamatsu-nanozoomer-med cleaned up profiles (again) no

hamamatsu-nanozoomer-med Align fiber on FITC -> 23% shading no

hamamatsu-nanozoomer-med Change fluo-bulb at 1934h no

leica-tcs-sp5-bio Hello, strange ticking noise when changing from one laser to another and changing takes a long time. It takes longer from blue to green than to red. Could not work. no

leica-tcs-sp5-bio Clicking noise from microscope and software getting stuck constantly. I think (not an expert) that something is wrong with the system that changes the filter cubes. The noise was coming from there and if I changed the filter cube in the SW the clicking noise started and the software froze. Did some images without changing the filter in between. no

leica-dm6b-navigator Training no

leica-tcs-sp5-bio Worked for 2 1/2 hours without problem. Suddenly a window opened "A motor device is jammed or has a general hardware problem (device id = Excitation FW UV (MA4-V1)!'' and the running stack could not end properly. After a moment it started again. Shut down the AF-Software, now it won't open again and tells, that it can't connect to scanner. no

visitron_csu-w1 Laser line measurement without objective: 405nm 0.3mW 445nm 0.98mW 488nm 2.8mW 514nm 0.25W 561nm 1.5mW

visitron_csu-w1 Fluctuation in laser power 488nm line are reported.

leica-dm6b-navigator Training no

hamamatsu-nanozoomer-med Training, no

hamamatsu-nanozoomer-med Laurence had today the following problem for the second time: The system got stuck in the Brightfield mode. Here the profile icons change from light bulbs to some blue id-card default icon. Also no profile is selectable (which makes kinda sense, given that in Brighfield mode there is not much that can be changed). Julien mentioned, that we have to make sure, that in the Settings, the 'Fluorecense Option' is checked. It was checked. What resolved the issue was restarting the NZ and the NDP.scan. So far we do not know which exact steps reproduce this problem. no

visitron_csu-w1 Training Nora Ruef (Stein Group) no

visitron_csu-w1 There is a problem with the connection between the new workstation and the EMCCD camera. It seems to be a problem with the plug to the board in the back of the new computer. We are working on the problem but for now if you lose connection to the EMCCD camera try the following: 1. Shut down VisiView software (only). 2. Unplug white camera cable in the back of the computer (see image attached). 3. Plug in cable again tightly. You should hear the computer making a sound, which indicates a new connection. 4. Start VisiView software. yes

leica-dm6b-navigator Updated LAS X to and display firmware to its latest version. The camera flipping problem is solved! Once it is set correctly the settings will remain persistent after a restart. DFC9000 -> horizontal flip. DFC5400 -> horizontal and vertical flip. yes

leica-tcs-sp5-bio Good morning, unfortunately there is neither lens cleaning tissue nor kintech paper. no

zeiss-axioplan-med The camera shows more and more hot pixels. On the right side of the chip the pixels seem to saturate quicker than on the left.

zeiss-axioplan-med Aligned the camera. But the stage movement still behaves erratically. The pitch seems to be different in different directions. Also when moving over long distances along the x or y axis, the shift seems to be oscillating.

visitron_csu-w1 Hey, I've got again the error message concerning the emccd camera. I've tried to move the white cable that I think is connecting the camera to the computer, I did not succeed in solving the issue. I could do my experiment with the cmos, so I did not spend to much time on it. Furthermore, I believe the red signal was unusually weak. Cheers, no

zeiss-axioplan-med Calibrations and Parafocality routines do not yield good results. The camera needs to be re-aligned and the slide holder needs to be checked if it has too much play. no

hamamatsu-nanozoomer-med The NDP.scan throws a the following error message when loading a profile. "input string was not in the correct format". To fix this: Control Panel > Region > Additional Settings > Numbers. The decimal symbol has to be a "." (dot) and the the digit grouping symbol nothing or a half-space. => Added a registry entry, that is loaded upon logon (c:hamamtsu\decimal-setting.reg). no

hamamatsu-nanozoomer-med the Cy3 has a date included in its name. This should be corrected via a remote session in the coming days. no

leica-dm6b-navigator Training: no

nanozoomer-s60 Network connection upgrade to 10GBit/sec. no

hamamatsu-nanozoomer-med Network connection upgrade to 10Bit/sec. no

hamamatsu-nanozoomer-med When selecting a profile the following error message shows up: "input string was not in the correct format". Workaround is to create a new profile. no

leica-tcs-sp5-med training no

nanozoomer-s60 Preventive maintenance: Cleaned and calibrated the instrument. The glue from the tape on the stage was hindering the loading process. The system is fully functional. no

nanozoomer-s60 Added the machine to the UNIFR domain. Please use your UNIFR credentials to login to the machine. no

hamamatsu-nanozoomer-med Cleaned up the user settings in NDP.scan. Please do not change the default profiles (first three items). Only use personal names for new profiles and use as few as possible. Alternatively you can use the NDP.viewer > Info functionality to figure out particular acquisition settings of a ndpi-file. yes

hamamatsu-nanozoomer-med Changed the computer for a new Win10 machine. User settings were transferred. The old machine will be kept as a reserve win7 machine for old hardware. no

hamamatsu-nanozoomer-med NDP.scan was updated to version 3.3.0. The old version is still there (no desktop shortcut though) since 3.3 is not officially supported for this instrument. no

hamamatsu-nanozoomer-med Preventive Maintenance: Cleaning of the instrument, alignment of the fluo-fiber (problem of overheating was not found, hopefully it will be fixed by the fact, that NDP.scan 3.3.0 will use the shutter of the lamphouse from now on). The stage causes vibrations and causes slightly blurry images, it will need to be replaced in the coming couple of months. Attenuation filter, that was put in manually for the calibration, is not needed anymore, since the software is not capable to avoid saturation itself - it was removed completely. Calibrations were ran, system is operational. no

nanozoomer-s60 What is the minimal maintenance doable by us? no

nanozoomer-s60 How to get the NDP.Viewer licence back from the old computer? no

hamamatsu-nanozoomer-med Check the status of the stage (noise reported by Julien Aupetit) no

nanozoomer-s60 If the software interrupts, or the state of the NZ requires a re-initialisation, the entered references are lost. no

hamamatsu-nanozoomer-med The light guide is leaking yet again. no

visitron_csu-w1 Windows backup does not complete with error 0x81000039. Try to run a backup on a USB HD. no

visitron_csu-w1 Added new win10 computer to the UNIFR domain. Added med.support_group and biol.support_group as well as to the Administrator group. no

visitron_csu-w1 Annual maintenance. [what was done?]

visitron_csu-w1 Changed win7 computer for a new win10 machine. no

zeiss-axioplan-med Updated to StereoInvestigator version 2019.1.3 no

zeiss-axioplan-med Changed the control computer to a new Win10 machine. no

leica-dm6b-navigator Training no

imaris-2 Installed 9.5 license file on the license server ( Uploaded the mac installer on Uploaded the windows installer on the Baramundi file server and made the necessary changes to the deployment script -> Imaris LATEST job now installs 9.5. no

leica-dm6b-navigator Extended Depth of Filed (EDF) post-processing does not work for tile-scans. Leica support did not provide a solution so far for the processing. However for EDF during acquisition does work with the corresponding software license (Trial license can be activated via help menu) no

leica-tcs-sp5-bio Hello, unfortunatlz the leica system does no work anzmore. It tells me: Hardware Server Update Error. May I just restart the programm? It told me to contact you. Thank you for your help. Franziska no

leica-mz-16-fa-med Training no

leica-dm6b-navigator Training: no

leica-tcs-sp5-bio There is no lens cleaning tissue anymore. Thank you. no

leica-tcs-sp5-med The port prism keeps getting stuck. When going live in the LASAF, there is no signal hitting the detectors. Aftter putting in the prism (top right button) and swtiching back and forth between binoculars and software the prism finally found its position again. (Forgot the sequence we found with Leo during hist last intervention)

leica-tcs-sp5-med Problem with scanning

leica-tcs-sp5-bio Training:, no

leica-tcs-spe-ii-bio channel 405 seems not to be detected anymore by the camera although it was very present through eyepieces. Cannot say for other channels.

leica-dm6b-navigator Service intervention by Urs Rüegg: update to LASX version 3.7.0. The bug for the DFC9000GT (monochrome) camera persisted. After each restart, the image needs to be re-oriented (toggle vertical checkbox). During my last use I even had the problem with the color camera, where it did not occur before. no

hamamatsu-nanozoomer-med Training no

hamamatsu-nanozoomer-med Cleaned objective (had some fiber at the edge of the field of view). Ran fluo. calibration: ok (cy3 and cy5 a bit low, about 80) no

ge-deltavision-elite-med Alain fixed the front door of the environment chamber. (The hinges broke during when dismounting the chamber for cleaning) no

leica-tcs-sp5-bio I could not use the confocal microscope Leica TCS SP5 (Bio) today. When I arrived everything was off and Windows gave me a message that it had not been properly turned off before. Then I went on and set up my experiment settings and when I tried to acquire it kept showing the message “Scan will start after target laser power is reached”, but it had been already one hour since I switched it on…Then I called you, went for lunch, came back, same message shows up...I doesn’t look like the laser is ever going to warm up. Since the microscope was off when I came, either the previous user didn’t use it or switched it off instead of leaving it in standby, so I tried to call to check if the problem was there before, but also couldn’t reach anybody. Have you seen that error before? Ana no

visitron_csu-w1 20x objective produces blurred images and needs inspection

leica-tcs-sp5-bio Training with Hendrik Oudhoff (Anna Jazwinska Group) no

leica-dmi6000b-bio Fast emission filter wheel was misaligned and loose (noisy). After consulting with A. Lüthi tightened the small screw on the side and cleaned the filters. no

ge-deltavision-elite-med training no

leica-tcs-sp5-bio Followup BSoD issue: After analysis of the c-drive, it turns out that it is very unstable, sometimes the drive can be mounted on another machine (icybox) sometimes not. Looking from a pmagic booted system, we find a pagefile.sys. The activated paging in Windows is probably what shortened the lifetime of the SSD. Cloning the old SSD on a new one did not work. Now I try to restore the windows system image from the last backup onto a new SSD. The backup is from the 26.08 which was just the day before Tomaz reported the error for the first time. The backup ran through. Creation of a repair disk was not possible. But a restore point was created today. Hopefully this will allow to repair the computer next time around. The pagefile size on C is kept to max 2GB. No Windows updates that where known to cause trouble (leica services) were found on this machine. no

leica-tcs-sp5-bio The computer does not boot. It stops in a Blue Screen of Death: Driver_IQRL_not_less_or_equal ... iaStorA.sys. no

leica-dm6b-navigator training, no

leica-tcs-sp5-med Replacement of the argon fibre. Alignment of the light path. The argon laser lines have now smaller than 50nm xy-shipts (UV/Vis), significantly more laser power output and the xz-psf profile of the 40x is more symmetric (still has a bit of a banana shape towards the left). no

ge-deltavision-elite-med Training no

ge-deltavision-elite-med The ventilator of the heating module is broken and is currently being replaced. No environmental control is possible until repairs will be finished. no

visitron_csu-w1 Training Mauro Sousa de Almeida (Group Barbara Rothen, AMI) no

leica-tcs-sp5-bio Training Flavia Fico (Ruegg lab) no

cosmos-workstation-med Moved computer from PER03 I.2.1 to PER02 0.410 yes

ge-deltavision-elite-med The field of view is again restricted, and this time already in BF (was only the case with fluorescent light when we fixed it with Boris on 22.07.2019). I've then check the camera settings that were problematic last time (Settings -> illumination -> filled stop open/close), but it was fine. I was not able to take any picture showing the restricted field, only had too much light or not enough no

threadripper Remote License Server (Imaris) was down, SysAdmin rebooted and now it is okey no

leica-tcs-sp5-med Wrong glass fibre was ordered, another order was placed by Leonardo Silva no

leica-tcs-spe-ii-bio Annual maintenance Leonardo Silva, system is calibrated and laser power values are above minimal values no

leica-dm6b-navigator script to copy 8/4 slide carrier settings does currently not work for new users. -> the Path variable was missing c:\windows\system32\ (with a tailing slash) no

ge-deltavision-elite-med Training no

leica-tcs-sp5-med Yearly maintenance: System largely ok. The transmission of the Argon fibre is bellow nominal. Laser coupling is not optimally adjustable, since the screws are already at the limit position. Fiber and coupling will need to be tested with respective replacement parts (quote ~12k). Consequently the PSF is a bit asymmetrical in Z-direction. no

leica-tcs-sp5-bio Yearly maintenance: PSF, Laser intensities, Chromatic alignment, galvo and scanner all have nominal values. System is in very good condition. (We need to consider buying a new computer for the Win10 update and the end of support of Win7) no

leica-dm6b-navigator Training no

visitron_csu-w1 Cannot import nd files with stiching info to OMERO

leica-tcs-sp5-med Display on the lamp house does not work anymore.

leica-dm6b-navigator Consulting: Ji Huang no

zeiss-axioplan-med Cleaned objectives, condenser and filters. Re-aligned the diaphragms and köhlered the system. no

leica-tcs-sp5-bio In between lines the UV laser power cannot be regulated for single scans. no

leica-tcs-sp5-bio Cockpit controls not responding, restart of software resolved problem no

leica-tcs-sp5-bio software crash after 2hrs scanning no

leica-tcs-sp5-bio Training: no

leica-tcs-sp5-med too small of the view of the objective no

leica-tcs-sp5-med Training;, no

leica-dm6b-navigator training:,, no

omero-2 Fixed: thumbnail regeneration of the histologie practicals. Logged in as histologie-de/histologie-fr, clicked on the datasets and waited until the thumbnaiis were created. There was a cumulation of things: During chown and changing group types to read-only the thumbnails were lost. Then there was a server space problem (it was full). Also see: no

leica-tcs-sp5-bio Training for Esther Martinez (Dengjel Group) no

leica-tcs-sp5-med training no

ge-deltavision-elite-med Training Simon Blanchoud (Blanchoud Group) no

ge-deltavision-elite-med Training Iga Tomczynska, Angel Martin Fernandez, Mohamed El-Shetehy (Mauch Group) and Riko Hatakeyama (De Virgilio Group) no

leica-tcs-sp5-med training no

hamamatsu-nanozoomer-med training no

leica-tcs-sp5-med training no

leica-tcs-sp5-bio Training Karin Wendrich (Albrecht Group) and Ana Humbert (Sprecher Group) no

leica-tcs-spe-ii-bio button Focus and Lower Z do not work. Strangely the same happened to Steve (at SP5 bio) one week ago. no

leica-tcs-sp5-bio Training no

visitron_csu-w1 No power on AOTF no

leica-dm6b-navigator Training Pierre-Yves Mantel no

leica-dm6b-navigator Prepare an empty slide for the shading corrections.

leica-dm6b-navigator Is there a way to save exposure settings for different objectives? Currently the exposure settings stay the same when changing objectives. -> Exposure is associated with the camera, not with the objective. no

leica-tcs-sp5-med Training + + no

leica-dmi6000b-bio the emission filter wheel fell out of position when switching fast. The issue was resolved by pulling the module out from the camera tube, removing the cover of the filter wheel an tightening the screw locking it on the axis no

leica-tcs-sp5-med Training no

leica-dm6b-navigator Camera splitter module exchanged, Replaced. Camera alignment done correctly. (by Urs Rüegg, Leica) no

leica-tcs-spe-ii-bio Training + no

leica-tcs-sp5-med Training no

leica-tcs-sp5-med Training no

hamamatsu-nanozoomer-med Could not use the fluorescence mode, only BF. no

leica-tcs-sp5-bio Restarted system three times and it is working again no

leica-tcs-sp5-bio lost signal mid-scan, emission to PMTs seems blocked no

visitron_csu-w1 Training Jianwen Zhou (Dengjel Group) no

leica-tcs-sp5-bio Cleaned the condenser module from glas debris. Cleaned the prisms. Köhlered the system. no

leica-tcs-sp5-bio Training Ora Hazak and Salves Cornelis (Hazak Group) no

leica-tcs-sp5-bio Broken glass was found in the condensor no

leica-tcs-sp5-med Depannage by Leonardo Silva from Leica: After checking the entire light path including the electronic components, Leo found that the camera port lense in the microscope stand got stuck. The scan head was removed and the components were cleaned. Then the entire system was aligned. no

leica-dm6b-navigator Narrowing down the image orientation problem: When the LASX is run as admin, this problem does not occur. Also since the update to 3.6.2 it is only the sCMOS (DFC9000) that alternates between vertical flips, The color cam (DFC5400) seems to be fine now. no

leica-dm6b-navigator Re-ran the LAS-X 3.6.2 installer. This time NVIDIA drivers were installed too. The firmware was not re-installed. no

leica-dm6b-navigator The orientation of the fluo images keeps flipping between sessions or even if we switch to the color camera and back. One time the image has to be flipped vertically, the other time no flipping is required. The annoying thing is that it has to be verified before each tile-scan. no

leica-dm6b-navigator When using the color camera, while being in 'Live' acquistion mode, switching between the color and b&w mode leads to all kind of color weirdness. Color settings need to be reset and white balance needs redoing. no

leica-dm6b-navigator Training no

leica-dm6b-navigator Updated to LASX 3.6.0. (Fixed shift between camera rectangle and the image on the canvas. New autofocus points) no

omero-2 Update the OMERO.insight clients to 5.4.10. updating Baramundi job (re-executed the job for all the machines it was already used). Put the executables (win, mac) on the no

leica-tcs-sp5-bio Training Marylene Bonvin (Jazwinska Group) and Brisselian Kamary (Sprecher Group) no

visitron_csu-w1 Training Guillermo Garcia Osuna (De Virgilio Group) no

leica-tcs-sp5-bio Training Steve Roubatel (Group Jazwinska) no

hamamatsu-nanozoomer-med Replaced light-guide of the fluorescent lamp. Ran successful fluorescence calibration. no

hamamatsu-nanozoomer-med The light guide of the fluo-lamp melted. Currently only BF imaging is possible. no

leica-dm6b-navigator training:, no

leica-dm6b-navigator training:,, no

leica-tcs-sp5-bio Changed condenser module. Configured 3 prisms. Koehlered. Adjusted the z-limits of the 20x. no

hamamatsu-nanozoomer-med Cleaned objective. It was smeared with glycerol. no

leica-dm6b-navigator training no

hamamatsu-nanozoomer-med Ran calibrations BF + FLUO -> ok Ran BF test scan -> ok no

leica-tcs-sp5-bio Confocal training no

hamamatsu-nanozoomer-med Cleaned objective. Some dust specs and some sticky stuff on the edge of the lense no

leica-tcs-sp5-bio The motor of the condenser lense does not seem to work. It would not move in when using high magnification objectives no

leica-tcs-sp5-bio Training Delphine Chinchilla (Group Weisskopf) no

visitron_csu-w1 Cleaned up c drive. Moved kaeserpe to d: (33GB). cleanup of windows updates. The content from 'c:\updates and drivers' I copied to \\bigdata\science\med\lmf\software\visiview (put a shortcut to the server in the folder). Its contents can probably be deleted after back ckecking with pablo no

leica-tcs-sp5-bio Cleaned up C drive (old users, windows updates cleanup) no

leica-tcs-spe-ii-bio Cleaned up C drive (was almost full) no

leica-dmi6000b-bio Gabriella Saro reported problems with USB connected Temp. Cherry Device. Maybe C drive needs clearing. no

visitron_csu-w1 Training for Rejane Carron (Group Weisskopf) no

omero-2 upgraded to OMERO 5.4.9 (server, on virtual-microscopy). Updated the client downloads. replaced the executables for the baramundi job and exectued the job for all the existing targets. no

hamamatsu-nanozoomer-med I did a BF scan and my scans were segmented and thus can not be used. The person after my had the same problem no

omero-2 Updated PathViewer to the latest pre-release. no

leica-tcs-sp5-med software can't open, I try to restart the computer but it still can't work. no

hamamatsu-nanozoomer-med There was shading rendering the scans unusable. There was some glue stuck to the objective. The objective was cleaned and the calibration re-run. The instrument runs normally again. no

leica-dm6b-navigator Leica Technician visit (Thibault). Could not fixate the tube for the cameras (reported the same problem with other clients). The pieces will be replaced. The cameras were aligned in the meantime. no

ge-deltavision-elite-med Training:, & no

hamamatsu-nanozoomer-med Preventative maintenance (Julien Aupetit): - General cleaning of the system and of the optics - Greasing of mechanical axes - Mechanical and optical adjustments - No software update, current versions: NDP.Scan 2.5.90, NDP.Calibration 2.4.9, NDP.View 2.6.17 - Brightfield bulb replacement - Calibration in fluorescence and brightfield - Tests on sample: 20x/40X, Brightfield/Fluorescence, Z-stack => OK Stage condition seems to start deteriorating. This can lead to some horizontal blurred stripes on the images, especially at 20x. A stage exchange will be recommended to recover perfect working conditions The system is working properly except this point. Replaced hallogen soket+bulb no

leica-tcs-sp5-bio Training & no

leica-tcs-sp5-med Reinstalled the 40x objective. no

leica-tcs-sp5-bio Training Rejane Carron (Weisskopf Group) no

leica-tcs-sp5-bio Training Sujana Sivapatham, Stefanie Wissmann (Stein Group) no

leica-tcs-sp5-bio Training Jun Abe, Lukas Altenburger, Juliana Barreto (Stein Group) no

leica-dm6b-navigator Bio-Formats MetadataL: OMERO cannot extract the channel wavelengths and the objective immersion. no

leica-dm6b-navigator Training & no

hamamatsu-nanozoomer-med Bright-Field calibration - ok no

hamamatsu-nanozoomer-med Fluorescence calibration - ok no

hamamatsu-nanozoomer-med Cleaned C: drive no

leica-dm6b-navigator Training, and no

hamamatsu-nanozoomer-med Changed the fluo lamp at 2007 hours. The new lamp has fissures in the ceramic holder of the lamp. Also the back of the lamp is yellowish instead of the usual nacreous color. Fluorscent calibration is still pending since there was an ongoing scan during the lamp exchange. no

leica-dm6b-navigator Training & no

leica-dm6b-navigator How to configure a multi-roi acquisition where the stage goes to a lower z-position when moving from one ROI to another?

omero-2 many connections on the server currently lead to a crash of omero.web, after which the nginx web-server will respond with a '500 internal server error'. The current workaround is a crontab installed as omero user, that checks every minute if the server responds with 500 and if it does it restarts omero.web. no

huygens-2 Install the beakerx kernel for the jupyterhub no

omero-2 after reboot of the, the omero server does not start automatically. Aditionally, there are two rogue processes: icgridregistry and glacier2router that need to be killed manually before the omero.server can be started again.

leica-dm6b-navigator Trainging & no

visitron_csu-w1 System has been upgraded and is functional again Laser output: 561nm 107mW and 55mW 515nm 61mW and 43mW 488nm 146mW and 115mW 445nm 90mW and 43mW 405nm confocal 59mW and 15mW 405nm FRAP 60mW and 35mW no

leica-tcs-sp5-bio Training Alexandra Siffert (Markus Geisler) no

leica-dm6b-navigator stage not aligned with slides positions (in the navigator), focus points set in focus map unfocus during image acquisition leading to blurred image. no

leica-tcs-sp5-bio Training & no

hamamatsu-nanozoomer-med Training no

leica-dm6b-navigator The tube for the filter wheel and the fluo camera, ist not well attached to the gate module. It has a little play that messes with the camera alignment. no

leica-dm6b-navigator Training and no

leica-dm6b-navigator the filter wheels do not have spectral data that can be seen in lasx. on mouse over it shows nothing. in the configuration tab I could not find the information yet. Either figure out to make the information visible in lasx or on the webpage

visitron_csu-w1 CSU-W1 unit and laser merge unit gone to Munich, for service and upgrade. Planned return on Tuesday October 23 no

visitron_csu-w1 Training Mohamed El-Shetehy (Mauch Group) and Dotun Adeleye Adeyinka (Egger Group) no

leica-dm6b-navigator Training and no

threadripper Installed 10GBit network card. Netmaster will put the plug 13.09 on the fast switch soon. Throughput should then reach about 0.5GByte/sec. no

leica-tcs-spe-ii-bio No JVM could be found on the system. Please define EXE4J_JAVA_HOME Needed for Omero no

leica-tcs-spe-ii-bio Stitching Tilescans does not produce good results with Convalaria Sample. Maybe needs calibration no

leica-tcs-spe-ii-bio 10x objective is dirty, probably full of oil, blurred vision no

leica-dm6b-navigator The fluo-camera (alreday) needs re-alignment. no

leica-tcs-sp5-bio during use the software cann't work no

leica-tcs-sp5-bio Training Dotun Adeleye Adeyinka (Egger Group) Virginie Trieu and Noemi Sgammeglia (Sprecher Group) no

leica-tcs-sp5-med The 40x 1.3NA was sent to Mannheim by Lüthi for an overhaul. The moving part of the objective is only actionable under considerable pressure. The objective needs to be taken apart and cleaned (attempts to clean from the outside have not been successful). The assymetry of the PSF will also be checked on this occasion. The overhaul will take 3 weeks approx. no

leica-tcs-sp5-med A. Lüthi managed to improve the alignment between blue and red. The 405 PSF is now well aligned, but particularly with the 40x 1.3NA, the PSF shows some assymetry, the scattering is more intense on one side of the Z axis. This intervention will not be billed seperatly but will be included in the maintenance contract (thanks) no

leica-tcs-sp5-med yearly preventive maintenance. The chromatic correction between blue and red had some problems. Maybe the system suffered slightly during transport. A. Lüthi will come back with the necessairy tools to try and improve the alignment no

leica-tcs-sp5-med Move from the basement to the new platform for fixed cells PER02 0.410 no

leica-tcs-spe-ii-bio yearly preventive maintenance. New Leica Tech. Leao was working on the microscpe and managed to improve the laser power at the objective lens. The LASX sofware was updated, hopefully this will take care of the metadata problem with the lif files (check). no

leica-tcs-spe-ii-bio Move from PER04 to the new platform for fixed samples in PER02 0.410 no

leica-tcs-sp5-bio yearly preventive maintenance. system was cleaned and aligned. no irregularities (laser power within specs) no

leica-dm6b-navigator Training no

threadripper Install the same network adapter as in the computer of the DM6B. Contact the netmaster to upgrade the connection (cable) no

leica-dm6b-navigator Training no

leica-tcs-sp5-bio 20x objective has dirt on the inside of the objective lense. Should be overhauled no

leica-dm6b-navigator Training with and no

hamamatsu-nanozoomer-med NZ training no

leica-tcs-sp5-bio C drive was full. Moved user data of users that had more than 1GB in their home folder to the D drive. PLEASE USE D:\ OR H:\ TO SAVE THE ACQUISITION DATA. yes

leica-dm6b-navigator Transmitted light source is broken. New one on the way. no

leica-dmrxa-med change the web-page address in the startup script (currently pointing to DMRBE)

zeiss-axioplan-med the instrument provoked a short circuit again. - repaired insulation on the cable connectors on the lamp house no

zeiss-axioplan-med Changed the fluorescence lamp at 585288 no

zeiss-axioplan-med The microscope pulled a short circuit when switching on the main switch. Identified the fluorescence lamp as having direct contact between phase and mass. Disassembled the lamphouse, shortened the copper wire on the cooler, measured resistances and re-mounted the lamphouse. no

visitron_csu-w1 Intensity values change over time in long-term movie for 561nm line no

visitron_csu-w1 Training Domenica Ippolito (Glauser Group) no

ge-deltavision-elite-med Training Domenica Ippolito (Glauser Group) no

leica-tcs-sp5-med confocal training no

leica-tcs-sp5-bio Mr Rodaro Alan from Leica is looking at the black screen no

leica-tcs-sp5-bio Monitor on microscope stays black no

visitron_csu-w1 Adjusted FRAP laser at end of fiber from 0.3mW to 3.1mW. It seems still low and needs to be experimentally tested. no

leica-tcs-sp5-bio Main switch PC/Microscope has been replace by Mr Andreas Lüthi no

ge-deltavision-elite-med the 40x OTF file is missing in the deconvolution menu.

leica-tcs-sp5-bio The green button PC/microscope can not be switched off - i.e. the microscope stays on. May it be switched off directly at it's power source? no

leica-tcs-sp5-bio left green button (PC /Microscope) cannot be switch off anymore no

leica-tcs-sp5-med Confocal training no

visitron_csu-w1 Training for Angel Martin Fernandez (Mauch Group) and Dina Muhieddine (Schneiter Group) no

leica-tcs-sp5-bio cleaned the computer, and worked around the bad contact of the back fans of the computer. The should be replaced asap. no

visitron_csu-w1 Installed Mosaic Macro/Plugin from Laurent Gelman (see also no

visitron_csu-w1 Training Maria Suarez Alonso (Hoogewijs Group) no

leica-tcs-sp5-bio Training Jiewen Zhou (Dengjel Group) and Martine Ferry (Puoti Group) no

visitron_csu-w1 488nm laser line again weak 405nm laser line weak, FRAP is not possible no

leica-tcs-sp5-bio Error message on screen at system start: 534-Rear Chassis fan(2) not detected no

huygens-2 The NVIDIA driver broke the remote desktop server XRDP (see So currently the only way to use the Huygens Essentials GUI is with X11

huygens-2 Updated Huygens Essentials and HuCore to version 18.04 no

huygens-2 Updated the HRM to version 3.5 no

huygens-2 Configured the HRM to use GPU (as admin user on the web, not in the config files) no

huygens-2 Installed nvidia drivers and CUDA libraries on the server to make the GPU usable by Huecore. no

leica-tcs-sp5-med The air-conditioning unit of the room seems to have been broken for a few weeks. The lights are on and the remote control works, but no cooling happens. It is extremely hot in the room (not good for lasers?) no

visitron_csu-w1 Laser fiber output (entry spinning disk) 405 2.5mW 445: 37 mW 488: 70 mW 515: 26 mW 561: 50 mW no

visitron_csu-w1 laser measurement on stage 20x objective: 405: 0.08 mW 445: 0.7 mW 488: 1.5 mW 515: 0.53 mW 561: 1.3 mW no

cosmos-workstation-med Hello, Imaris doe not show the arena sections anymore. This section contains all samples so it' is not possible to poperly use the software for batch analysis. I was still fine 10 days ago and it is the same version 9.1.2 I used and clicked on the licence for batch analysis. Thanks for your help Nils no

hamamatsu-nanozoomer-med Bright-Field calibration. OK. no

hamamatsu-nanozoomer-med Removed the glass debris from the loader, the NZ is scanning normally again. no

hamamatsu-nanozoomer-med The slide scanner throws an error message concerning the wrong slide format after loading the first slide. The scan is immediately after interrupted. Restart of the instrument is necessary thereafter. OUT OF ORDER until further notice. no

ge-deltavision-elite-med Replaced the multiple socket in the rack. The switch on the rack multi-socked melted yet again. So far I was unable to find a switch that has the proper circuit breaker properties to avoid the electric arcs. no

visitron_csu-w1 Error message for calibration of ablation laser Wrong live acquisition setting no

visitron_csu-w1 488nm laser line is weak, about 50% output based on standard sample no

visitron_csu-w1 Training Flora Raez (Schneiter Group) no

ge-deltavision-elite-med the general switch number 1 is not working. I've used the other switches to turn off the light sources. no

leica-tcs-sp5-bio No more cleaning tissues no

leica-tcs-sp5-bio the lever of the gate for the transmitted light PMT is loose again. no

leica-tcs-sp5-bio confocal training no

ge-deltavision-elite-med training: no

leica-tcs-sp5-bio Training Stien De Coninck, Corine Dos Santos Reis (Rüegg Group) Training Aboubakr Moradi (Mauch Group), Valeria Zoni, Stefano Vanni (Vanni Group) no

hamamatsu-nanozoomer-med Booking rules: During office hours maximum 4h per day and per person. If you need to book for longer periods, you can book over night or over the weekend. The goal is to allow to use the instrument by as many people as possible (also with small batches of slides) no

visitron_csu-w1 Visiview upgrade to version 4.1. New functionality: Tilescan module, python macros. Please test your macros. if there is any trouble please take screenshots of the error messages and send them to us. (see yes

leica-tcs-sp5-med Fixed the SSD tray and replaced the previously mentioned temporary solution with a definite solution. Long live the SSD! no

leica-tcs-sp5-bio confocal training no

leica-tcs-sp5-bio the scanner produces visual patterns (horizontal and vertical lines). Need to talk to the service technician about it no

leica-tcs-sp5-bio The wheel console (gain, offset, pinhole, etc) did not react - probably was not initialized properly. no

leica-tcs-sp5-med Exchanged main hard drive (WD HDD) to a Samsung EVO PRO SSD. Cloning went without trouble. Extended the main partition to fill the entire SSD (now 512GB). The tray for the drive was adapted temporairly. It needs to be slotted properly for long term. no

leica-tcs-sp5-med Swapped the memory bars (slot 0 and slot 2), dusted the hardware pieces. Now, the computer booted normally. We need to keep an eye on this. Memory can be changed by us. The computer replacement is recommended to order from Leica. no

leica-tcs-sp5-med confocal training: no

leica-tcs-sp5-med Memory test failed during boot of the computer. Probably need to replace the DIMM RAM soon no

leica-tcs-sp5-med Generally slow processing, long login times, memory error message, restart was necessary , reported by Desirée König no

leica-tcs-sp5-med The green laser or argon laser does not work or is too weak. Please can you cancel my reservation as I could not take any images. no

leica-tcs-spe-ii-bio Gwenn Jaquier Renaud (Egger Group) no

leica-tcs-sp5-bio Training Gwenn Jaquier Renaud (Egger Group), Stephanie Käser-Pebernard and Christine Vionnet (Dengjel Group), Cyrielle Kaltenrieder (Sprecher Group) no

omero-2 Update Baramundi job to 5.4.3. Launching installation on the bioimage computers. no

omero-2 updated to OMERO 5.4.3 no

leica-tcs-sp5-med training: no

leica-tcs-sp5-med Training: no

hamamatsu-nanozoomer-med fluo-calibration: ok no

imaris-2 Put the 9.1.2 dmg-file (Mac) on the download page. Updated the Baramundi Imaris job and ran it on the Image processing workstations. no

imaris-2 Update RLM license files. We have now 4 Base, Measurement, Vantage and XT. no

threadripper could not start Imaris on the machine. Imaris was not responding. Generally slow processing? no

leica-tcs-sp5-bio confocal training: Katya Nardou no

leica-tcs-sp5-bio Compile User List for Access to PER17 013 no

visitron_csu-w1 Training with Shadi Abou-Eid (Dengjel Group) no

visitron_csu-w1 Training with Aboubakr Moradi (Mauch Group) no

leica-tcs-sp5-bio Regine Toelle reported Error Controller V1 after an hour of scanning. Problem reoccured after restart. Subsequent users did not report problems. Interogated with Mr Lüthi at Leica. Response: Lieber Herr Egger ZB Bei einem Neustart, wenn man etwas zu schnell mit Staren der SW war, hatte ich schon ähnliche Fehler. Wenn es während dem Betrieb auftritt, so ist es möglich, dass sich der eine oder andere Schrittmotor verschluckt und die Kommunikation sich dabei verhedert hat. Wenn es nicht regelmässig auftritt und ev noch eine Zusatzmeldung wie Stepper Motor 1….oder ähnlich heraus gibt, so ist es schwierig etwas genaueres darüber zu sagen. Meist hängt es aber mit dem Scanner zusammen, wenn das System sich auf diese Art verschluckt. Ich würds erst mal weiter beobachten. Wenn es beim Aufstarten passiert ist, so würde ich auch noch die Revolverstellung am Mikroskop prüfen. Wenn da die Position nicht eingerastet ist, so gibt es auch komische Fehler heraus, weil das Abfragezeitfenster überschritten wird. Ä schöne Tag und Grüsse Andreas Lüthi no

visitron_csu-w1 Ablationlaser shows uncontrolled movements in x axis. Remote Sesssion with Visitron and switching xy-cable indicates fault in Galvo controller box. Sent Galvo controller to Visitron via FEDEX. Fault was repaired by Visitron and Laser Ablation system is working again. no

leica-tcs-sp5-bio Curtains are installed at new location PER17 013 no

visitron_csu-w1 New Immersion Oil bottle arrived Nikon 50 Type N MXA22166 index=1.518 no

leica-tcs-sp5-bio Organise date for curtain installation no

visitron_csu-w1 Adrien reported problems with ablation laser. Check on Monday no

omero-2 Upgraded PathViewer to 2.1.0. It now allows markdown syntax in the description (see, hence we can have section titles, links etc. no

imaris-2 Imaris 9.1.0: Put the installer for mac on the download page. Created Baramundi job and installed on the numbercrunchers. no

threadripper installed Imaris, Matlab, Matlab RC, Citavi, Endnote, Photoshop, Illustrator, Anaconda, no

imaris-2 Updated license files on the RLM server for Imaris 9.1.0 no

threadripper The screen menu is flickering on and of unprompted. Benq service was contacted. no

threadripper Inventorized the machine, installed LAN network connection and joined UNIFR.CH domain. From now on users need to use their UniFr credentials to log on to the computer. no

visitron_csu-w1 Note the computer inventory number on the maintenance page and set the variable for medpc43 in baramundi no

leica-tcs-sp5-med The 40x 1.3NA objective only moves with difficulty. We will have to send it to Germany for a complete overhaul. This will cost ~1/4 of the purchase price (~3k) no

leica-tcs-sp5-med Annual service: cleaning and calibration no

leica-tcs-spe-ii-bio Check the metadata of the files after the latest maintenance and software update (before the metadata import in OMERO for this system was not working properly)

leica-tcs-spe-ii-bio Annual service: objective and condenser cleaning. Software update LASAF 3.1.2. Köhler Illumination configs. The 10x was used with oil, which was hard to clean. no

leica-tcs-sp5-bio Replacement of the Argon Laser and power supply. Shortening the ventilation tube for more cooling efficacy. The lever for the transmitted light gate PMT was loose. no

visitron_csu-w1 removed the 32GB hyperx ram set. For some reason it broke the computer initialization. no

hamamatsu-nanozoomer-med introduction no

hamamatsu-nanozoomer-med Did a test with a new light guide. After calibrating both light guides, there was no siginificant difference, neither in intensity, nor in patterning. no

hamamatsu-nanozoomer-med The illumination patterns might be connected to the old optical fiber (mentioned by a technitian that does nanozoomer service in the U.S.A). We should discuss this with julien and consider a replacement at the latest for next service. no

zeiss-axioplan-med 10x glyc. is broken, the lens does not move anymore. yes

zeiss-axioplan-med cleaned condenser, filters and binoculars. Aligned diaphragms. no

zeiss-axioplan-med changed fluo-lamp. counter: 569953 no

imaris-2 Updated to 9.0.2. Created baramundi job and run it on the ip workstations. put the mac installer on the download site. no

visitron_csu-w1 Upgrade: swapped the 1TB HDD with a 2TB SSD. The old HDD is still in the computer, but not connected to the main board anymore. no

visitron_csu-w1 Upgrade: added 32GB RAM, now it's 64 in total (the maximum the main board can handle). The Ram configuration is not quite optimal, but good enough. no

leica-tcs-sp5-med Currently, when switching on the live mode, the shutter closes automatically after a couple of sections, while the scanning still continues. This black images. To work around the problem, the live-button needs to be clicked two times, once to stop the scanning, and then to recommence it. no

leica-tcs-spe-ii-bio Frequent error messages, restart required no

leica-tcs-sp5-med confocal intro: and no

leica-tcs-sp5-med Confocal Intro: no

leica-tcs-sp5-bio Argon Laser seems to be dead. Enduring problems with ventilation in PER17 might significantly shortened lifetime . no

leica-tcs-spe-ii-bio Microscope Error: DM 5500 Controller Class:1 Code:15 Text: cmserial2:FT_IO_ERROR (errorcode:4) Please restart system I had to restart three times today with this error message no

leica-tcs-sp5-bio air extraction not working. No 488nm laser visible no

huygens-2 Moved the HRM form svx-uo7640hrm2 to srx-bioimage, the new 'real' server. no

leica-tcs-sp5-med confocal intro: no

leica-dmi6000b-bio Issue nr1: sometimes the camera view is different from the eyepieces view and I see part of the field that is very bright and part of it black. If I restart the software and the microscope by the controller, usually the issue stops (but the next time I use the microscope it happens again). Issue nr 2: during time lapse recording, the background intensity increases in one frame and comes back to normality in the following frame. This happens anytime during a long recording (1h), disturbing the measurements because I get a noisy trace. Again, sometimes it happens, sometimes not. (I have some examples to explain this more easily) no

omero-2 upgraded to OMERO.server 5.3.5 (zeroc-ice 3.6, postgres 9.6, java 8) no

omero-2 upgraded server to debian 9 (preparing to match the new HRM server) no

omero-2 Fix the maintenance redirect and page when omero web is down. Currently it shows only a nginx 502 "bad gateway" error. no

leica-tcs-sp5-bio Error report server connection lost. Screen freezes. Two incidents today with Julien Perler. Look at error log. no

leica-tcs-sp5-bio Training Gabriella Saro and Filipe Marques (Glauser Group) Training Lana Rees (Jazwinska Group) no

leica-tcs-sp5-med Conforcal Intro:, no

hamamatsu-nanozoomer-med Nanozoomer intro no

huygens-2 When renewing the server certificates with Switch include the new DNS: and the new alias: no

huygens-2 put a place holder in the username field on the login page of the HRM. no

leica-tcs-spe-ii-bio User training Florent Aziri (Egger group) no

leica-tcs-sp5-bio User training Clarisse Brunet and Jenifer Kaldun (Sprecher lab) no

leica-tcs-sp5-bio Curtain fittings need adjustments in PER17 013 no

ge-deltavision-elite-med Make a ghost image (backup) once the AD configs and OMERO installation are done no

ge-deltavision-elite-med Introduction:,,, no

ge-deltavision-elite-med The main switch (white) creates big electric arcs. It is a matter of time before this (new) switch too will melt. The switch needs to be replaced with a better switch. no

ge-deltavision-elite-med Login with Active Directory is currently not working. Use the "worx" and serial number as credentials as workaround. yes

leica-tcs-sp5-med confocal intro: no

leica-tcs-sp5-bio ventilation not working and thus argon laser not working today at 9 o'clock. Problem with ventilation should be solved now (BE) no

imaris-2 Got the Imaris 9.0.1 (bugfixes). Adapted the Baramundi job and launched for biolpc94, biolpc79, medpc211 and medp213. For mac the installer is on the download page. no

leica-tcs-sp5-med confocal intro no

leica-tcs-sp5-bio Problems with Argon laser shut downs. Probably caused by interruption in the air extraction system (400m3/hr). Bimetallswitch secures overheating of the laser and switches off the power. Restart is necessary to reinitiate the laser. Sanitär engineer and Siemens are working on the problem. no

leica-tcs-sp5-bio Hello, I have had a few, what I think are, software problems since the confocal moved. The first is a problem with the set up of tile scans. After I mark the positions and I start the imaging or go back to live, the positions shifted. It is only a little, but really annoying when I position my sample and after I made the picture the top of the sample is just of screen. The second problem might be related is that the merge often is poorly done (which I can fix myself with Fiji). The last one is that during the merge the program stops working regularly. You then have to restart the program, re-initialize, etc. I hope my message makes sense Stage was not screwed on properly. Problem solved (BE) no

ge-deltavision-elite-med Move from PER09 to PER05 no

leica-tcs-sp5-bio Ask for LAN connection plug no

leica-tcs-sp5-bio Maintenance: Calibration of the system. Change the slide holder's attach point. Argon and Helium 543 Laser should be observed, possible replacement maybe next year. Define focus positions to use the "focus" button and "lower z" button to quickly move the stage. no

leica-tcs-sp5-bio Umzug von PER05 to PER17 no

leica-tcs-sp5-bio Training for Julien Perler (Jazwinska Group) no

visitron_csu-w1 cleaned filters in CSU-W1 with Pablo Radermacher no

cosmos-workstation-med installed Imaris 9.0 no

imaris-2 Installed 9.0 license no

leica-dmrxa-med Adjusted, koehler illumination, objective configurations, and limit positions for all objectives no

leica-dmrxa-med The buttons for the z-axis (table up and down) are not working properly, both of the buttons make the stage move up, not respecting the limit positions. This can crash the slide. Use the wheel with a big step size to move the table down. yes

leica-tcs-sp5-med Confocal intro and no

hamamatsu-nanozoomer-med There exists a update for the NDP.scan which simplify the acquisition parameter tweaking (without having to leave the scan). We should consider updating (Helmut will make an offer) no

hamamatsu-nanozoomer-med Maintenance: Change the fluo-lamp. Given that it had only 189h, we will get a replacement. The light guide was rearranged, to be able to change the lamp without touching it. The system was cleaned and realigned. no

hamamatsu-nanozoomer-med The fluorescent calibration for the DAPI did not run through. The system complains about the light source alignment. Remote session Julien Siue for a remote session to do the alignment is scheduled for week 31. no

leica-tcs-sp5-med confocal intro no

leica-tcs-sp5-med Confocal intro Michelle Kueffer no

leica-tcs-sp5-bio Move of TCS Leica SP5 to new building September 5-6, 2017 Contact: no

ge-deltavision-elite-med Move to PER05 planned for September 6-8, 2017 Contact: no

visitron_csu-w1 20x objective is cleaned and focussing is restored no

leica-tcs-spe-ii-bio Cleaned up the c: drive. no

visitron_csu-w1 Unable to focus with 20x objective. Possibly contaminated with oil. Needs cleaning. no

leica-tcs-sp5-bio Old User profiles should be deleted. C: Harddrive is full. no

visitron_csu-w1 User Jie Liu could not access sCMOS camera while other User had no problem. no

visitron_csu-w1 Filters are quite dusty, flat field correction is needed for Brightfield images. Filter needs cleaning soon.

leica-tcs-sp5-med The computer does not boot. Blue-screen appears, but not long enough to get the error message. The microscope is not usable until this problem is solved. no

hamamatsu-nanozoomer-med Introduction with no

hamamatsu-nanozoomer-med Ran fluorescence calibration. no

hamamatsu-nanozoomer-med Changed fluorescence bulb at 1110 hours. The counter was reset. The bulb showed some small bubble in the filament. no

visitron_csu-w1 The file nameing seems to be off. Sometimes there is an underscore to separate bits of information and sometimes not. no

visitron_csu-w1 The instrument meta-data is incomplete. Visitron Systems should configure the software during the next maintenance to write ALL the paramter into the metadata file. no

visitron_csu-w1 Check the x-y absolute position in the metadata of the images. Currently we observed that PositionX and PositionY is always the same for all images that are part of a tiling acquistion.

imaris-2 The RLM (floating license server) is now hosted on a new VM ( This was necessary to update the RLM to v11.3. For the users nothing should change. The new VM has the same alias as the new one ( See IT > Linux Servers > bioimage-srv for more details). no

omero-2 Installed no

omero-2 Updated OMERO.figure to 3.0.0. no

omero-2 Update the 5.3.1. This fixes the isssue with the public data access. Currently the public user is automatically logged in for the image viewer (^webclient/\d{3,10}/) no

imaris-2 Unable to checkout licenses. The RLM license server got hung up (yet again). Fixed by reboot of the VM. no

leica-tcs-sp5-bio Training Jie Liu (Geisler Group) no

leica-tcs-sp5-bio Training Lucia De Andres Bragado (Sprecher Group) no

omero-2 Upgrade OMERO.server to 5.3.0 no

huygens-2 Updated the HRM to 3.4.1. no

hamamatsu-nanozoomer-med Currently it happened several times that scans of entire brain sections showed a pattern of stripes (top down) due to non-homogeneous illumination. A fix was proposed by Hamamatsu and we executed it. Please write an incident for this instrument in case you observe these patterns. no

leica-tcs-sp5-med Confocal intro: (Filgueira Group) no

leica-tcs-sp5-med How to import spectra into LASAF for probes that are not in the software catalog yet? no

leica-tcs-sp5-med The Hoechst emission spectrum is wrong. The instrument configurations need to be adjusted. no

ge-deltavision-elite-med The signal for the CFP/DAPI is generally really low... maybe the lamp should be exchanged? no

ge-deltavision-elite-med Introduction for (Schneiter Group) no

ge-deltavision-elite-med Introcution session with and for marco capogrosso's group no

ge-deltavision-elite-med The keyboard driver is not properly installed. The numeric keypad does not work (e.g. 0 is insert)

ge-deltavision-elite-med The main switch (white one) that should also control the camera is not working. no

ge-deltavision-elite-med The DAPI configuration when using the binoculars lets through a lot of other light. I have been unable do a proper sample positioning on the nuclei. The FITC and mCherry work though. no

ge-deltavision-elite-med The 40x oil OTF file for the deconvolution seems to have gone missing. The PSF has to be reacquired.

leica-dmi6000b-bio Port needs to be activated to connect computer to network no

visitron_csu-w1 Ablation Laser is not working, We see the laser dot when we want to calibrate but is too low to calibrate. Furthermore, the laser is completely unfocused and move alone. no

leica-tcs-spe-ii-bio Images taken with 532nm laser line show horizontal stripe pattern. Other laser lines are fine. In contact with Leica Technical Service Team. no

leica-tcs-sp5-bio 633nm laser line drops in intensity mid-acquisition. In contact with Leica Technical Service Team. no

leica-tcs-sp5-bio Training Zsofia Csaky (Schneiter Group) no

leica-tcs-spe-ii-bio Training Zsofia Csaky (Schneiter Group) no

leica-dmi6000b-bio Leica DMI6000 has been physically moved to the live cell platform PER05 0.1.301b (Basement) no

leica-tcs-sp5-bio While using the button LOWER Z, I noticed the microscope makes a click-click sound. I did not notice this before. Also, when using the touch panel to select 40x, it got stuck during rotation. This did not repeat when trying it again. no

leica-tcs-sp5-med confocal introduction: no

leica-tcs-sp5-med The button to switch between objectives of different immersion types does not work (microscope stand) no

leica-tcs-sp5-bio Short Arc Lamp changed and counter reset. no

leica-tcs-spe-ii-bio Leica annual service 2016, Mathias Fournier no

leica-tcs-sp5-bio Leica annual service 2016, Mathias Fournier no

visitron_csu-w1 Training Zsofia Csaky (Schneiter Group) no

imaris The RLM license server executable delivered by Bitplane are not 32bit compatible anymore (even though we initially installed the 32 bit system because they only had 32 bit). We'll need to upgrade the bioimage-srv to a 64 bit system at some point. no

imaris the license server was unresponsive (web-interface) and did not checkout licenses to clients. Had to reboot the VM. no

leica-tcs-sp5-bio Training with Jules Duruz (Sprecher Group) no

leica-tcs-sp5-bio Training with Simon Blanchoud (Jazwinska Group) no

huygens Images can be transferred from OMERO to HRM but not always the other way around. When transferring data from HRM to OMERO, and the target dataset belongs to a group that is not currently set as the users default, then there will be an omero.SecurityViolation. The current workaround is to set the users default group accordingly (see no

visitron_csu-w1 Laser alignment. Input into CSU-W1 was doubled for all the wave lengths. 45 min work time will be billed. no

visitron_csu-w1 488nm laser line is weak and needs reajustment no

leica-tcs-sp5-med Confocal maintenance: Calibrations and alignments done. The stage drift on beads was not reproducible. There was a decrease in certain of the argon wavelenghts (458). Given the current utilization the argon laser will be probably due for replacement in a year+. no

leica-tcs-sp5-med Confocal intro: no

leica-tcs-sp5-med start up protocol is not the same as on the website no

leica-tcs-sp5-med There are no Lens Cleaning Tissues in confocal room no

leica-dm6000b-med The light intensity regulation for the fluorescence is not working. After startup the attenuation wheel reacts on the shutter button. Then it takes the position of the max intensity (simple hole). After this the wheel reacts once only to the first click on the intensity buttons and then remains blocked. Not sure if it is the buttons (bad contacts or the servo motor) no

leica-dm6000b-med The servo-motor of the condenser lens is not working. Also the lever between the motor and the lens is broken. Exchange ~2k no

leica-dm6000b-med Changed and aligned Fluo-Lamp. Counter: 30676. Koehlered and cleaned the microscope (a lot of salt on the stage, seeping into the condenser) no

leica-tcs-sp5-bio Training Tomaz Martini and Iwona Olejniczak (Albrecht Group) no

visitron_csu-w1 405nm laser has been re-coupled into fiber for FRAP application. 12mW measured at the end of the fiber. no

visitron_csu-w1 New GFP filters with narrower bandwidth were installed. Suitable for combination of GFP and Nile Red. 510/20 no

imaris The license server did not respond. During the startup of Imaris, to software would hang just after showing the splash-screen. The server was rebooted. The reinitialisation of the RLM seemed to have fixed the problem. no

visitron_csu-w1 405nm laser cannot be activated. Restart of entire system sometimes fixes problems but laser seem weak. FRAP application does not work. no

huygens Database connection error. Blocked the web-interface on the landing page. no

leica-tcs-spe-ii-bio The metadata of the lif-files is incomplete. The microscope configurations should probably be adapted in order that this information get's written to the experiment file too. no

leica-tcs-sp5-bio Training Magali Jungo (Sprecher group) and Regine Tolle (Dengjel group) no

hamamatsu-nanozoomer-med Fluorescence Calibration. OK. no

hamamatsu-nanozoomer-med Replace fluorescence lamp. Operating time: 1769 no

leica-tcs-sp5-med when I switsched onn the Resonance scanner a highfrequencz noise starteed no

digital-slidebox There is no SSL!

digital-slidebox The translations in German and French are bad to the point that they are ambiguous.

digital-slidebox The request new password functionality doe not work. The error message reads: "There was a problem with sending the mail."

digital-slidebox the user registration only works with the english language. In french it fails

zeiss-axioplan-med updated StereoInvestigator to 11.09 no

hamamatsu-nanozoomer-med Maintenance (Julien Siue). Cleaning of the instrument, repair of the door screw, alignment and calibration of the light path and detectors. no

leica-tcs-sp5-med Confocal Intro: Smart & Flavien no

leica-tcs-sp5-med The 3D viewer in ImageJ does not work at this time. no

leica-tcs-sp5-med intensity values oscillate in the live mode when keeping all the parameter stable (stage postition, gain, etc.) no

leica-tcs-sp5-bio Occular/Camera Port Shutter is stucked midway. Confocal Scanning is still functional. Leica Engineer is informed. no

leica-tcs-sp5-med Replacement of the TL-Switch module. Cleaning of the optical parts (condenser, lens on the PMT). Alignment and adjustment of the mirror for homogeneous illumination. no

visitron_csu-w1 Training Lucia Lichvarova (Beat Schwaller Group) no

leica-tcs-sp5-bio Training Pengchao Hao (Markus Geisler Group) no

leica-tcs-sp5-med some users reported some weird drifts during their acquisition. It's not sure if it is imprecise movement or stage drift. Bead-acquisitions have to be made to investigate. no

visitron_csu-w1 Power interruption to change circuit board for entire PER05 no

visitron_csu-w1 Live imaging chamber set up with Joel Bourquin no

leica-tcs-sp5-med confocal introduction no

imaris Imaris 8.3: put the mac installers on the bioimage download page. Created Baramundi job for the windows machines. no

imaris Updated the RLM license server with the 8.3 licenses. The software will be released at the end of May. no

leica-tcs-sp5-med confocal intro no

leica-dmi6000b-bio Training Dedy Septiadi (Barbara Rothen Group) no

leica-tcs-spe-ii-bio Training Filipe Marques (Glauser Group) no

leica-tcs-sp5-med Changed Fluorescence Lamp. no

huygens ims-files can't be passed from hrm to omero no

huygens OME HRM bridge: the selection for the source image or the target project/dataset is not recognised. HRM complains: "Please select a destination dataset within the OMERO data tree." no

zeiss-axioplan-med updated StereoInvestigator to 11.08 no

ge-deltavision-elite-med installed the OMERO client. Making a shortcut in the application menu would not work. Users who register for the first time should have a desktop shortcut. For existing users the desktop shortcut was copied to their home folder (couldn't test it though) no

leica-tcs-sp5-med Today was not possible to use fluoerescence light to observe sample via eye. After pressing the incident light shutter buttom (open/close) there was no fluorescence light. The only possibility how to observe sample was to focus on sample in bright field. no

visitron_csu-w1 Training Joel Bourquin, Dedy Septiadi, Daniel Hauser (Group Barbara Rothen) no

huygens There were trouble with lif files. ImageJ would open them but the HRM job failed. Turs out it was a double-space in the file name, that would cause an error on the HRM no

imaris Got the 8.2.1 update. There are a couple of bug fixes. Unfortunately the confusion between the time and z-dimension hasn't been fixed yet. The Baramundi job was created and tested on the number cruncher. for mac the image is available on the download page (as usual) no

leica-dmi6000b-bio enlarged the c: partition. now it has double the size. before there were only 8GB left with all user data cleaned out. this wouldn't have been enough to handle the user accounts. no

leica-dmi6000b-bio installed firefox, java, imagej, omero, endpoint protection (anti virus) and Baramundi no

leica-dmi6000b-bio install some 127 windows updates no

leica-dmi6000b-bio installed regular backup with the MedAdmin user (saved on \\bigdata\science\med\lmf\backups\dmi6000b-bio) no

leica-dmi6000b-bio Hooked the computer up to the UniFr domain no

leica-tcs-sp5-bio Since this week there are problems with mosaic merge. Photo's are not stitched together nicely. no

hamamatsu-nanozoomer-med Another slide got stuck during the loading procedure. Since cariage was not reatreaded from the loading rack, the user was unable to open the loading door. When continuing the slide was then broken. The cause was probably rough slide edges (overlapping cover lid or label). no

leica-tcs-sp5-med The servo-motor of the mirror switching the transmitted light between the ocular and the PMT is broken. Current the mirror is in the position so the transmitted light can be used on the microscope itself (finding the sample position) but no transmitted light images can be acquired. no

leica-tcs-sp5-med Confocal Introduction Flavia Fico no

leica-tcs-sp5-med Filter wheel got stuck. No light on the eye piece or in the LASAF software. no

huygens certain images cannot be transferred with from OMERO. HRM complains that it can't find a given image ID (even though it exists when checking with OMERO.insight) no

huygens Jobs hang with status paused in the queue without ever being processed no

huygens Updated to HRM version 3.3. Changed the VM in order to update to Debian 8.3 and allow to install ice-zeroc-3.5 for the OMERO bridge. no

hamamatsu-nanozoomer-med A slide fell down (and ist still inside the instrument). It has to be removed (and probably the slide holder needs some cleaning). no

omero Updated to 5.2.2. Changed the VM since an OS upgrade was necessairy. Updated the clients on the download page and the baramundi jobs. installed the clients on the microscope computers. no

leica-tcs-sp5-med Replacement of the optical fiber of the UV channel. Laser power at objective plane: 0.8mW (before), 6mW (after). no

leica-tcs-sp5-med The mirror gating the bright-field to the transmitted light PMT or the eye-piece is stuck. So bright-field can be acquired via software, but when working on the microscope itself, the mirrors shadow is obscuring the sample. no

leica-tcs-sp5-med Confocal Intro: Lichvarova no

leica-tcs-sp5-med Introduction for Smart and Diego no

leica-tcs-sp5-bio User Training with Sivlia Almeida (Sprecher Lab) no

leica-tcs-spe-ii-bio Upon starting up the system, the optical path to binoculars is always blocked. Only upon restart of the entire system shutter opens. Suspect that laser safety shutter is faulty. Discuss with Mathias Furnier this Thursday no

huygens Submitted jobs to online deconvolution are not processed. Even 24 hours after submission. no

visitron_csu-w1 with mCherry channel settings there is dirt in the light path. Probably on the filter again in the CSU-W1 no

leica-tcs-sp5-bio There seems to be a problem with mosaic acquisition and stiching. Accuracy of stage movement needs to be assessed no

ge-deltavision-elite-med Sometimes the driver for the network adapter is not loaded properly, with the result that computer has no network connection. So if you cannot login because the system says that there are not login severs, please restart the computer. yes

ge-deltavision-elite-med The number keys of the keyboard do not work. NumLock has no effect. no

hamamatsu-nanozoomer-med There is a problem with the storage of the pictures in the batch mode / BF+ F. Only 1-2 of 3 pictures are stored randomly out of the three needed (BF, Cy2 and Cy3). no

ge-deltavision-elite-med The centralized switch (white) is currently not working. The microscope controller, computer and camera (switch on the adapter bellow the computer) have to be switched off individually. no

huygens The admin email of HRM needs to be changed no

omero The update of java to version 8 seems to have broken the OMERO clients. no

cosmos-workstation-med disk space upgrade: boot partition 2 TB SSD, Work space (buffer) 8 TB. Memory upgrade: 64 GB. The old data is made available on the drive W:_old_accounts\ yes

hamamatsu-nanozoomer-med fluorescence calibration run through successfully no

hamamatsu-nanozoomer-med The lamp controller came back repaired. The replacement was send back to Hamamatsu. The counter of the original lamp controller is still at 372h and is now operated with the lamp we kept as spare during the time of the repair. The lamp that busted was exchanged for free against a new lamp (new type). Also see incident 248 no

leica-tcs-spe-ii-bio The computer is now connected to the intranet and part of the UniFr domain. Use your university account to login to computer. In case you need to access your old local account, use .\ before the username, (for example; .\Administrator) and the corresponding password. no

leica-tcs-sp5-med First: when I arrived at 7:30, (3) (microscope stage) and 5 (laser fan) were still on. When I switched on all the other buttons, I only received error messages: no valid licence to start scan, CM-stick (last error code=200). Impossible to start LAS AF. no

leica-tcs-spe-ii-bio inventorized the computer: 38660, no

huygens updated to 15.10 no

leica-dm6000b-med Always me reporting issues! ;) I wanted to change from eye-mode to camera-mode, but this "screw" (?) that I would draw out (on top left) for this purpose, has lost its internal contact... I tried to "screw it in" again but did not manage to do so --> Taking pictures is impossible at the moment! (Diana) no

leica-tcs-sp5-bio Training Session Darbon Geoffrey (Group Didier Reinhardt) no

ge-deltavision-elite-med computer doesn't boot no

ge-deltavision-elite-med Software update: JRE to Java 1.7. (In case the java version needs to be restored, just link /usr/bin/java/ -> /usr/java/jdk1.5.0_11/bin/java) no

leica-tcs-sp5-med confocal intro no

leica-tcs-sp5-med Confocal introduction Janine Woerthmueller no

leica-tcs-sp5-med Confocal Intro Diana Roccaro no

leica-tcs-sp5-med Confocal introduction Luis Filgeira no

leica-tcs-sp5-med Introduction Jan Marro. no

visitron_csu-w1 Computer cannot start. Ask if I want to start using restoration system. Should I? If not I don't know how to turn off the microscope no

hamamatsu-nanozoomer-med The closing screw of the door is missing its blockage (so it does not fall off). This should be fixed during the next maintenance. no

hamamatsu-nanozoomer-med Replacement of the lamp controller. old lamp had 372 hours. This lamp currently serves as reserve. The new controller came with its own lamp and 110 hours on the counter. Our controller with the busted lamp was sent to Hamamatsu for investigation. Fluorescence calibration was run through successfully. no

leica-tcs-sp5-bio There is still a software bug in LAS for setting parameters (emission window etc) for channels in a new user account. Will talk to Mathias Furnier about this. Boris no

leica-tcs-sp5-bio Two users reported that when switching from 10x to 20x the objective crashed into glass slide. There is some bug in parafocality settings. I reprogrammed the settings and it works in my account, however I am afraid it may happen again for other Users. Informed Mathias Furnier and David Uhlmann. Boris no

leica-tcs-spe-ii-bio I today booked slot from 9-12 (morning) and I got confirmation through e-mail. However my slot was not displayed and Boris was able to overwrite it. Can u please sort out this issue for me. no

leica-tcs-sp5-med Installation of the 40x/1.3NA and correction lenses for the new objective as well as the correction lenses for the 20x/0.7NA no

leica-tcs-sp5-med 20x obj dirty with immersion liquid causing blurry images. Cleaned-up. no

visitron_csu-w1 Training Martina Korber (Schneiter Group) no

visitron_csu-w1 Issue with Perfect Focus, digital display says PFS DIS, PFS Focus wheel does not work no

visitron_csu-w1 LED light, I did not manage to adjust Green component and UV component. Blue component was fine. no

hamamatsu-nanozoomer-med Replaced Fluo-Lamp at 859h. no

leica-tcs-sp5-bio Ar Laser is replaced and fully functional Laser output: 138 mW Fiber output: 120 mW on objective plane: 6.0 mW no

leica-tcs-sp5-med confocal introduction: and no

leica-tcs-sp5-med confocal introduction with Boreggo Ines no

leica-tcs-spe-ii-bio Training with Lola Hostettler (Glauser Group) no

leica-tcs-sp5-bio Training with Lola Hostettler (Glauser Group) no

visitron_csu-w1 Pablo Radermacher (Visitron) repaired laser safety, aligned laser lines and cleaned dichroics in CSU-W1 unit. Drop in 561nm laser might be caused by faulted laser safety circuit. Currently, all laser lines are fine. no

visitron_csu-w1 Training Session of Daniel Hauser (Group Barbara Rothen) no

leica-tcs-sp5-bio Ar laser has very low output. Ar laser output 32 mW Ar fiber output 6mW on objective plane: 0.26 mW Acquisition of new Ar laser is in process. no

visitron_csu-w1 Drop in 561nm laser power reported. Visitron Engineer is coming next Thursday 16.07.2015 no

leica-tcs-sp5-bio Leica Engineer is coming this Friday afternoon to adjust laser lines. no

ge-deltavision-elite-med Main power switch is broken. Leave it as it is (in active position) and use power button on individual components (see picture) for the time being. no

leica-tcs-sp5-bio Across a number of confocal sessions, the most recent being 01.07.15, I have experienced a drop in (perceived) Argon laser power. Using the same laser power under the "Configuration" and "Acquire" tabs, I noticed that after ~30 minutes the intensity of both the fluorescent and brightfield channels dropped noticeably. After this drop, the intensity either fluctuates briefly between low and high, before staying low, or stays low. no

ge-deltavision-elite-med From COTTIER Stephanie: Ai fait une session de microscopie ce matin sur le Delta vision. Quand j'ai voulu éteindre la console centrale (qui commande l'allumage de l'ordinateur), rien ne s'est produit, les indicateurs lumineux verts sont restés allumés. no

visitron_csu-w1 Please be aware that currently the automatic laser safety shutter is not working. Hence never look trough the eyepieces or look directly at your sample when you are in live or acquisition mode. For safety close the occular shutter before going to live mode. Before you open occular shutter and look trough occular make sure lasers are not active. If you have questions don’t hesitate to contact me or any other member of the Bioimage Team. no

visitron_csu-w1 Training with Rasha Khaddaj (Schneiter Group) no

visitron_csu-w1 Laser security shutter closed in middle of acquisition for a few frames. Other Users reported similar problem with shutter. Will check with Pablo Radermacher (Visitron). no

leica-tcs-sp5-bio Mr Ruegg from Leica has installed an new touch screen. All the buttons on microscope are on the previous configuration. no

leica-tcs-spe-ii-bio The computer is not yet part of the unifr domain no

leica-tcs-spe-ii-bio The microscope configuration seems to be out of date. Some metadata is incorrect. The leica specialist should check this. no

leica-tcs-sp5-bio Meeting with Leica Engineer Rüegg. Sent wrong piece to repair Touch Screen. Next Meeting Friday June 19th. no

leica-tcs-sp5-bio Training with Pauline Fritsch (Sprecher Group) no

leica-tcs-sp5-med confocal introduction no

leica-tcs-sp5-med there were repeatedly crashes with the software when exporting images to tiff. no

leica-tcs-sp5-bio Touch Screen does not respond no

huygens the following user cannot save any templates: no

leica-tcs-sp5-med Manish Jain confocal session. no

zeiss-axioplan-med put a new ethanol "pissette" and some lense cleaning paper. no

leica-tcs-spe-ii-bio Maelle Bochud (Jacob Group) no

leica-tcs-sp5-bio Training Session Thomas Vorley (Metraux Group) no

visitron_csu-w1 Training Session Riko Hatakeyama (De Virgilio Group) no

visitron_csu-w1 Training Session Bejamin Pillet (Kressler Group) no

hamamatsu-nanozoomer-med To avoid unnecessary long acquisition times please use, depending on the ROI size use 3-10 focus point. Since the focus plane is fitted to these points, more usually does not result in better focus. For uneven samples use smaller ROI's. (avoid 5 points due to software but) yes

hamamatsu-nanozoomer-med Please use to inform your colleagues of notable booking-deviations, so they can profit from the time not used. yes

leica-tcs-sp5-med Confocal plane is not parallel to object whatever the objective. Pb with stage or light path alignment? no

leica-dm6000b-med Problems with Camera settings since some days; contrast is insufficient so that e.g. 4x4 binning has to be used (or Rolling average / ...) in order to (well) see a signal (I just wanted to try out the Report Issue function; Christoph knows already...) :) no

leica-tcs-spe-ii-bio Hard disk is full... It affects the software function. could you ask users to empty their folders? Thank you no

leica-dmrxa-med Köhlered and aligned the microscope. no

visitron_csu-w1 No signal for GFP (both cameras) althoug under the microscope it can be seen that GFP is highly expressed no

leica-dm6000b-med Installed Hokawo software no

leica-dm6000b-med The computer was had to be completely re-installed (disk and OHCI card were damaged). Please use the HCImageLive software for image acquisition. The Hokawo software is being procured. no

leica-tcs-sp5-med changed glycerol bottle no

visitron_csu-w1 Visit Pablo Radermacher, Fixation of FRAP module, Calibratrion of FRAP and Ablation Laser, 1.5x lense Macro. Decided to clean filters in CSU-W1 once Plexiglasbox is constructed no

leica-tcs-spe-ii-bio User training session Gaia Bonetti (Egger Group, Biology) no

leica-tcs-spe-ii-bio User training session Dimitri Vanhecke and David Thevenaz (Barbara Rothen, AMI) no

leica-dm6000b-med The computer does not start. Consequently no images can be acquired currently. A blue-screen is shown with the message: The registry cannot load the hive (file): \SustemRoot\System32\Config\SOFTWARE... no

hamamatsu-nanozoomer-med Replaced the lamp at 574h (the replacement was sent free of charge). See incident 190. no

hamamatsu-nanozoomer-med The fluo-lamp broke. Only bright-field imaging is possible until further notice. no

hamamatsu-nanozoomer-med The instrument has focusing problems. no

leica-tcs-sp5-med introduction Parejo Begona no

leica-tcs-sp5-med Confocal intro no

visitron_csu-w1 sCMOS camera gives blinking view (reported by no

leica-tcs-sp5-med Confocal introduction Jérôme Widmer and Ines Mottas no

leica-dm6000b-med Dear users, There is a problem with the epifluo. The y axis doesn't work so in not possible to move the slide. Did someone use the microscope today? Thank you, have a nice evening. Manu no

leica-tcs-sp5-bio Received an error message at 12:30 when starting a live scan: Hardware Server Update Error Microscope error: DM6000 Controller Class: 1 Code: -1 Text: cmserial2:FT_IO_ERROR(errorcode:4)d Then, while closing the LAS program: Runtime error! Program:D:\ProgramFil... This application has requested the runtime to terminate it in an ususual way. Please contact the application's support team for more information. no

leica-tcs-sp5-med Manish Jain confocal consulting no

visitron_csu-w1 Again dust on filters that is visible with 100x and long exposure time no

leica-tcs-sp5-bio There is no lense cleaning tissue left - thanks! no

leica-tcs-sp5-bio The settings for the light of the bright field were changed and don't fit with a normal usage. no

leica-tcs-sp5-med Introduction with Michael and Solange no

leica-dm6000b-med Fluo introduction for no

hamamatsu-nanozoomer-med Replaced fluorescence lamp at 1028h. Reset of the operation hour counter. no

leica-tcs-spe-ii-bio Training Maryam Syred (Sprecher Group) no

leica-tcs-sp5-med Confocal session Manish Jain no

leica-tcs-sp5-med Confocal session with Manish Jain no

hamamatsu-nanozoomer-med The fluorescence lamp does not start correctly. Consequently the Nanozoomer is usable only in bright-field mode. no

leica-tcs-sp5-med modified the stickers on the microscope for the filter buttons to fit the current configuration no

leica-tcs-sp5-med Exchanged the stage controller no

leica-tcs-sp5-med Assisting Manish Jain during the confocal session. no

visitron_csu-w1 Overlay image is fixed in upper left corner. no

visitron_csu-w1 Cannot always hear shutter noise when laser safety is turned on. no

visitron_csu-w1 EMCCD camera lost connection: PM1394 Cam00: Camera communication problem, please check connections! Restart of entire system reinitialised connection no

leica-tcs-sp5-bio In order to acquire old user settings go to: G/ Documents and Settings/"Your User Name"/Application Data/LeicaMicrosystems/LAS AF/Instruments Paramers Setting Copy Settings to: C/Users/"Your User Name"/AppData/ Roaming/LeicaMicrosystems/LASAF/Instrument Parameters no

leica-tcs-sp5-bio Ar laser initially did not power on. Had to restart entire system. no

ge-deltavision-elite-med Introduction for no

visitron_csu-w1 Ablation laser needs again recalibration. Looks like FRAP module is not stable positioned. no

visitron_csu-w1 Maintenance: Visitron Kaiser and Radermacher serviced system and recalibrated FRAP and Ablation module no

visitron_csu-w1 Training Chi Zhang (Scheffold Group) and Alessandro Sardu (De Virgilio Group) no

leica-tcs-sp5-bio Maintenance (Lüthi & Fournier). The microscope was thoroughly cleaned, especially the condenser (there was quite a bit of oil there). The control computer was replaced. AOTFs were calibrated. Ar Laser line is weak, probably needs replacement 2015 no

leica-tcs-sp5-med Maintenance (Lüthi & Fournier). The lasers were double checked. The light path for the bright field that had some shadows on the lower magnification objectives was realigned. no

leica-tcs-sp5-bio The control computer is now part of the UNIF.CH domain. Please use your university account to log onto this machine from now on. no

leica-tcs-sp5-bio x-y control wheels for the stage are not independent. When I turn the X wheel, the stage moves both in X and Y direction, and inversely. no

leica-tcs-sp5-bio Before computer change, control panel (smart gain, smart offset,...) did not always connect correctly. It turned on and I could spin the buttons, but the numbers wouldn't change. To resolve the problem one needed to switch off the PC, switch off the PC/microscope button and switch everything back on. I don't know if the problem will still occure now that PC was changed. no

leica-tcs-sp5-med The 63x gly objective was not cleaned at all when I came in the afternoon - full of glycerol. no

leica-tcs-sp5-med Light intensity is unevenly distributed across the field in the SP5 of Medicine. This started weeks ago, but now it is more annoying. When removing one of the oculars, it can be seen that there is something misaligned which produces this black shadow. Javier no

visitron_csu-w1 Installation of system by Visitron (Peter Kaiser and Pablo Radermacher) no

visitron_csu-w1 Demonstration ONIX Microfluidics Station Group Schneiter and Group DeVirgilio no

leica-tcs-spe-ii-bio Training Jan Marro and David Pury (Jazwinska Group) and Larisa Georgiana (Sprecher Group) no

leica-tcs-sp5-med Confocal introduction Travis no

leica-tcs-sp5-med The button to switch between imersion and air objective does not seem to work (no reaction) no

visitron_csu-w1 Training Dimitri Vanhecke and Dagmar Kuhn (Barbara Rothen) no

visitron_csu-w1 Lightpath Widefield UV is not correctly configured, there seems to be a shutter closed. no

leica-tcs-sp5-med confocal introduction with wassim eid. no

visitron_csu-w1 Training Shirish Mishra , Stephanie Cottier, Vendula Stradavola, Magdalena Marek (Schneiter Group) no

ge-deltavision-elite-med The servers LMF, Common and Home should be mounted. (contact Pawel after 19.Oct.) no

leica-tcs-sp5-med During gain adjustments the laser power shuts down after short time (DPSS561 laser). One can hear clear click. no

visitron_csu-w1 No light path settings for FRAP or Ablation with 100x no

visitron_csu-w1 FRAP and Ablation laser needs calibration no

visitron_csu-w1 Köhler Illumination for 100x can't be achieved since condenser position is limited. Need to adjust entire condenser position. no

visitron_csu-w1 Dirt in light path need to be traced. no

visitron_csu-w1 Training Marie-Pierre Peli-Guli, Marta Moreno Tores, Malika Jaquenoud (Group De Virgilio) no

visitron_csu-w1 Training Oriane Guillermin (Egger/Sprecher Group) no

leica-tcs-sp5-med Maintenance: Microscope was cleaned and the lasers aligned. AOTF and AOBF have been calibrated. Not all issues have been adressed (132, 133) they will be once the technician comes back for the bio confocals. no

leica-tcs-sp5-med The 405 and 655 Lasers seemed to weaken. Their power is still in the specs, but they might need replacing soon. no

leica-tcs-sp5-med The optical fiber seems to have bad transmission and might to be changed soon. no

ge-deltavision-elite-med There are some samba problems. Files cannot be copied to a server from the linux machine no

ge-deltavision-elite-med There is no objective cleaning stuff no

ge-deltavision-elite-med There is no CO2 for the life cell chamber no

ge-deltavision-elite-med The laser module is currently not aligned properly, so it does not hit the marked roi's on the image. no

leica-tcs-sp5-med There seems to be a stage drift (reported by Franziska). After long observation times (20min) the position changes. Needs to be tested no

visitron_csu-w1 Training Adrien Vaquie (Jacob lab) and Lena von Giesen (Sprecher lab) no

leica-tcs-spe-ii-bio Introduction Shahnaz Lone (Sprecher lab) no

leica-tcs-sp5-bio Introduction Travis Carney (Sprecher lab) no

leica-tcs-sp5-bio Workstation Leica SP5 Bio does not boot. Leica Engineer Mr Fournier is informed and we evaluate options no

leica-tcs-sp5-med confocal introduction for manish, jain no

leica-tcs-sp5-med XY precise stage controller does not works properly no

leica-tcs-sp5-med As reported also by Manu, the UV laser lost quite a bit of intensity (probably dis-aligmnent). So you might need to work with increased UV-Laser power, as compared to usual. no

leica-tcs-sp5-med Confocal introduction for BERNARDO GARCIA Francisco Javier no

leica-tcs-sp5-med The button on the right side of the microscopes to switch the objectives does not allow to select all the objectives. It needs to be reconfigured. no

leica-tcs-sp5-med There is a shadow in the bright field image. When taking out the eye piece, there is something in the light path I could not identify yet. Possibly some diaphragm no

leica-tcs-sp5-med The stickers for the filter cube buttons need to be adjusted to the information on the website. no

leica-tcs-sp5-bio The wheel of the stage controller (bottom one) is loose and needs tightening no

leica-tcs-sp5-med The oil for object 40 x is almost finished and the bottle is a little broken because outside of the bottle is too slippery and fell down on the floor. I am so sorry for my carefulness. However, there is not leakage. It's better change a new bottle. Thank you very much. no

leica-tcs-sp5-med When I switched on the confocal on 11th at 8h, the bright field from the microscope looked like someone before wanted to adjust bright field settings. Only the center was lightened. I don't know how to put the bright field back to normal conditions. So I let it like this. no

leica-tcs-sp5-bio Introduction Vendula Stradalova (Schneiter Group) no

leica-tcs-spe-ii-bio Please close immersion oil bottle after use! no

leica-tcs-sp5-med Dear users, the whole system was left ON overnight. Please try to avoid such incidents in the future! no

leica-dm6000b-med Changed Mercur Lamp Counter: 22186 no

leica-tcs-sp5-bio Error message when starting LAS AF software: blocked DIC Turret. This means that you need to lower the condenser so that the stage can be initialized. Boris yes

leica-tcs-sp5-bio I booked the confocal this morning. Unfortunately when I started the software, it was indicated something like "Error with the Turret", and I could not work. Nothing changed even after restarting the sofware or the computer. The error had been detected by two users from yesterday. Boris easily identified and solved the problem. It was a problem with the DIC turret. B. Chassot no

leica-tcs-spe-ii-bio Introduction Makhabbat Saudenova no

leica-tcs-sp5-bio Introduction Makhabbat Saudenova and Isabelle Spuehler no

zeiss-axioplan-med The firewire card was exchanged to the one recommened by Hamamatsu. The system is operational again. no

leica-tcs-spe-ii-bio Oculars, objectives and microscope stand cleaned. no

zeiss-axioplan-med The firewire card delivered by the DIT appears not to be compatible with the Orca 05G camera. A new card is ordered and will be installed a.s.a.p. no

leica-tcs-sp5-med confocal introduction no

leica-tcs-sp5-med Confocal introduction. Suresh and Alina. no

leica-tcs-sp5-med Since the last maintenance the UV laser power dropped considerably. So you might need to user a bit more laser power on the UV than usual. no

leica-tcs-sp5-med Cleaned the objective and the eye piece lenses. The 40x had a lot of oil, it started to flow down in the objective wheel. Cleaned the stage. Measured laser power UV at 100% - 40x: 48uW - 20x 181uW - 10x 222uw no

leica-tcs-sp5-med I didn't get good images for DAPI with the 40x objective. Quite often there is this problem using the 40x. no

leica-tcs-sp5-bio New Leica Glycerin Immersion has arrived. Please use Type G Leica Immersion liquid from now on. no

leica-tcs-sp5-bio Laser ventilation was on over night, microscope not covered. Error Report: Cannot connect to hardware server. After restart everything was fine. no

hamamatsu-nanozoomer-med Please report all incidents when a slide falls down (Instruments>Slide Scanner>Hamamatsu NanoZoomer (med)>Report Issue). To get to the bottom of this we need know the frequency of such events. If you happen to sit at the microscope, please also note the procedural step (loading un-loading, etc.) the sound and place where the slide has landed. Thanks for your collaboration no

zeiss-axioplan-med Replacement of the control computer. no

leica-tcs-sp5-bio Introduction Adrien Vaquie no

leica-dm6000b-med Cleaned and köhlered the microscope. The right eye-piece has a fissure. The condenser lense mechanism is not working (for a long time now) no

leica-tcs-sp5-med Confocal introduction. Jelena Zaric no

leica-tcs-sp5-med Dear Felix, I was not able to use the confocal tonight. Franziska used it before me and started having issue with the Galvo. She restarted the program and there was a run time error appearing. Can u look into this and delete my booking on 12.05.14 from 7-8pm? THX& have a good night, LA no

hamamatsu-nanozoomer-med There were a couple of slides falling down during unloading. The system (maybe not optimally calibrated) seems to be more sensitive to minor irregularities on the slides. PLEASE CHECK CAREFULLY THE BORDERS OF YOUR SLIDES (IDEALLY BEFORE MOUNTING YOUR SAMPLE) AND REMOVE ANY GLUE, OR OTHER OBSTACLES. no

hamamatsu-nanozoomer-med The slowness of the NDP Viewer is due to the Microsoft security essentials. The process needs to be excluded from surveilence to work around the problem. no

hamamatsu-nanozoomer-med The arrangement of the fiber and the swtich for the fluorescence lamp is not obtimal. When pushing the switch on and off one always risks to touch the fiber and compromise the alignment. The fiber should be attached better. Think about taking the switch from the lamphouse and put it with a cable away from the fiber. no

hamamatsu-nanozoomer-med Maintenance/Repair Session no

leica-tcs-spe-ii-bio Introduction to Emmanuelle Muenger no

hamamatsu-nanozoomer-med Cleaned the slide-holder of the instrument and two successive batch scans completed successfully again. (see incident 80) no

leica-tcs-sp5-bio Since the Leica Engineer updated the pinhole and software we had incidents of software crashes/freezing. Please report any such future incident in the Report Issue Form. Boris no

hamamatsu-nanozoomer-med The slides are falling down when unloading from the scan-stage. They need to be removed by opening the scanner (Franck or Felix). DO NOT USE UNTIL FURTHER NOTICE. no

leica-tcs-sp5-med Confocal introduction to no

hamamatsu-nanozoomer-med The door locking mechanism is currently birdged, so it is possible to open the door while the NanoZoomer is scanning. DO NOT OPEN THE DOOR WHILE THE ORANGE STATUS LIGHT IS ON (BUSY)! no

leica-tcs-sp5-med The pinhole was replaced. The LASAF software was updated. The UV-Laser was a bit fine-tuned (increase from 1, 2 to 2.3mW). no

hamamatsu-nanozoomer-med The instrument has currently problem with the slide loading (solenoid-error). The Nanozoomer is not usable at this moment. (The Hamamatsu technicians are informed). no

leica-mz-16-fa-med The 1x objective has some small cracks

cosmos-workstation-med Introduction to Imaris and some general information about image processing to: Christof Wyss no

leica-tcs-sp5-bio installed two new computer screens no

leica-tcs-sp5-med Laser power button ON again this morning. no

leica-tcs-sp5-med Introduction for Adrien Vaquie no

leica-tcs-sp5-bio SP5 Bio will be serviced by Leica Thu 20/3. no

leica-tcs-sp5-med SP5 Med will be serviced by Leica Thu afternoon and Fri 20/3-21/3. no

leica-tcs-sp5-med The second time this week (after Monday morning), the LASER POWER BUTTON WAS ON OVERNIGHT AGAIN. no

leica-tcs-sp5-med Hello, I could not get the DIC picture. I can see the cells in BF in the microscope but when I switch to computer I could not see them. Sincerely, Shirish Mishra(8660) no

leica-tcs-sp5-med I booked the SP5 confocal in Med Dpt but I didn't manage to see my sample at the microscope with the bright field. The person before me didn't use the bright field. So I don't know if it's a trouble in the settings or if the lamp should be replaced by a new one. My collegue will need also the bright field this thursday. Could you let me know if she should book another confocal? Thanks. Laure Bapaume no

leica-tcs-sp5-med I would like to draw ur attention on the fact that the last 2 times (19.02, 03.03)we have booked the confocal. the DAPI was not working properly. On this base, we were not able to take paper quality picture and would appreciate if the fee would be waived.I am not sure why the UV light is still unstable...have u had any other reports? they have the pictures and we can send/show them to you. THX for your help, Lavinia no

leica-tcs-sp5-med Dear Confocal users, recently we found that the laser power button wasn't switched off, probably meaning that it was on all the weekend... Please, try to avoid such incidents in order to maintain the best performance of the lasers! no

hamamatsu-nanozoomer-med Dear Nanozoomer users, if there is no more storage space on the computer of the NanoZoomer, it will become impossible to make new scans. Please, move and then delete all your data which is still on the PC to free some space! Thanks in advance, Viktoria no

leica-tcs-sp5-bio Introduction with Adrien Vaquie (Jacob Group) no

cosmos-workstation-med Berenice Chassot: Introduction Huygens, Imaris no

leica-tcs-sp5-med Confocal intoduction Nils Lannes. no

leica-tcs-sp5-bio Introduction with Geisler Group no

leica-tcs-sp5-bio FlexScan computer screen is showing vertical stripe pattern and is flickering. For the time being I have exchanged the screen with a smaller one, which works fine. Boris no

leica-tcs-sp5-med The lasers have been on overnight. Apparently, the last user forgot to switch off or to announce that he/she will not come.... AGAIN!! no

leica-dm6000b-med The Demo R2 camera was sent back and replaced by our old C4742-95-12NR. The tests in Japan did not show any sign of malfunction, thus no problem was confirmed by Hamamatsu. After installation the image acquisition works again with the Hokawo software no

leica-dm6000b-med The Hokagawo software does not manage to capture an image at this time. The HCimage works without problems. no

leica-dm6000b-med Köhlered the microscope and cleaned the condenser lense no

leica-tcs-sp5-bio I booked the confocal this afternoon. Unfortunately when I started the software, it was indicated "Fatal error: no board detected", and I cannot work. Nothing changed even after restarting the sofware or the computer. Laure Bapaume no

leica-tcs-sp5-med The technician from Manheim, Hefner, was here. He re-calibrated the pinhole and was not able to reproduce the irregular behaviour (inconsistency between the signal and shown pinhole opening) after that. According to him the pinhole might not have to be changed after all. The alignment of the UV laser was fine-tuned (more power gets though now). The new shutter for the argon laser was installed. no

leica-tcs-sp5-bio Markus Hefner Leica Microsystems re-calibrated Pinhole. There was no need to exchange pinhole. Guarantee work after software update to LAS AF 2.7.3 no

leica-dm6000b-med The microscope need adjusting no

leica-tcs-spe-ii-bio Claire Jacob reports that the software is frozen. Restarted PC and software is working again. no

leica-tcs-spe-ii-bio Anne-Sophie used the SPE II this afternoon and reports: My session went quite well. I got much better quality pictures than before. I was doing mosaic and stacks at the same time. It crashed the software 3x. It's probably a little too much for the software… Here is the error message that I got: Please close software Unexpected error on scan DM55oo stage Ctrl Class:1 Code:-1 Text: amserial2:FT – 10 – ERROR (errorcode:4) no

zeiss-axioplan-med Re installed the LEP Stage no

hamamatsu-nanozoomer-med Changed the mercury lamp. Counter: 1748 hours no

leica-tcs-sp5-med The DAPI signal appears to be generally weak. The light path of the UV has to be checked and fine-tuned. no

leica-tcs-sp5-med Maintenance Session with Andreas Lüthi: The laser was realigned and The AOBF was changed (Incidents 20 + 37). The shutter did not arrive in time for installation. But Leica will offer a new pinhole, since apparently the weird behaviour is caused by a bad firmware. Hence we will get the pinhole on warranty. The session lasted form 9:15-14:45 no

leica-tcs-sp5-med Hi Felix, This morning I found the confocal on, apparently the user who booked it for yesterday afternoon didn't come or didn't switch it off. We wrote already thousands of times that people should always come when they book, especially if it's the last user of the day. But it continues to happen again and again... Could you please write a round email to the users of the group to remind them the user policies? Thanks a lot, Viktoria no

leica-tcs-sp5-med The shutter of the Argon laser needs to be replaced. Careful, tilting backwards the condenser of the microscope does currently not interrupt the light path of the Argon laser. Make sure to switch of "live mode" before manipulating your slide! no

leica-dm6000b-med New Cy5 filter cube (equivalent from Chroma) for far red imaging. no

leica-tcs-sp5-med UV laser not available this week. no

leica-tcs-sp5-med The Argon laser doesn't work - again. There seems to be a problem in the light path of this LASER, which could not be identified yet. We are getting the Technician form Leica to look into this asap. no

cosmos-workstation-med The Imaris License server is not working. Hence Imaris is currently not usable no

leica-dm6000b-med Installed Demo Camera ORCA R2 from Hamamatsu no

leica-dm6000b-med Camera can't communicate with the control computer. no

leica-tcs-sp5-med Changed the Glycerin bottle no

leica-tcs-sp5-med The pinhole behaves a bit weirdly. (it starts beeing shut around 80 microns already). Seems to be similar to the problem we already had. Also we should change the unit of the pinhole parameter into airy-units instead of microns. no

zeiss-axioplan-med The Microscope is not in use at this time. The stage has been sent to MBF to resolve the precision errors we were recently experiencing. The Stage should be back in about two weeks (23. Oct.) no

hamamatsu-nanozoomer-med There is a problem with the mechanic arm or the stage, the slides fell down. no

leica-tcs-sp5-med Increasing the overlap of the frames til 25% (instead of 10%), didn't solve the issue. The mosaic was still improperly reconstituted. no

leica-dmrxa-med Installed the Windows 7 32 bit system so that VisiView can interface with the camera again (there is no API for third party software to interface with the camera driver) no

leica-tcs-sp5-med Einführung für Silke Lehman und Studenten no

zeiss-axioplan-med The stage has an impresision when changing direction in x-direction. it's off about 7um. The Instrument is workable, but you should undertake a backlash correction before using it. no

cosmos-workstation-med The machine can't boot anymore (blue-screen) no

zeiss-axioplan-med The Stage makes a strange noise when reaching the end of travel in x-direction. Workaround: try to avoid travelling all the way to the end of x-axis. no

leica-tcs-sp5-med The sliders for the PMT detectors seem to have crosstalk if they are close together. no

leica-tcs-sp5-med Lüthi Maintenance session no

leica-tcs-sp5-med For the AOBS1 Leica borrowed us a driver. Our driver just broke and needs to be repaired or we have to buy a new one. no

leica-tcs-sp5-med 20x objective replaced. The new objective is a confocal-grade, dry one with better corrections and higher NA (0.7). no

leica-tcs-sp5-med Replacing the ventilators of the AOTF visual and AOBS1. no

leica-tcs-sp5-med The control computer is now part of the UNIFR domain. From now on all the confocal users have to login with their university login. If you need to login as TCS-User: medpc255\TCS-User and use the usual password no

cosmos-workstation-med Installed FlowJo on the machine no

zeiss-axioplan-med Stage faulty: sometimes does not respond when trying to move in the x axis. Can be temporarily fixed by restarting control box. Problem with either the joystick or the controller. Please check with MBF (Braul) and give fedback to Federica. no

leica-tcs-sp5-med The images acquired with bright-field illumination show a pattern (shade on the left third of the image) no

leica_tcs-sp5-med 20x/Imm (multi-immersion) broken, replaced by 20x/0.5 Dry (Lent by Leica). no

leica-tcs-sp5-med Changed Glycerin immersion liquid bottle. (it was empty far too soon) no

leica-dmrxa-med Get a Spot camera adapter and make the configurations for micromanager to control the camera no

leica-dmrxa-med Upgrading the control computer. The new machine drives and controls the Spot color mosaic camera with the Spot software. no

cosmos-workstation-med The R-server crashes if the input table contains a Bitmask type (image processing plugin). Since with R there is no manipulation to be done on such data, the workaround is to filter the columns. On the other hand the server needs to be restarted. no

leica-dm6000b-med Replacing screen with 24'' display. no

leica-dm6000b-med Rearranged the working place. Installed new table, Rollconatiner and chair. no

leica-tcs-sp5-med Get new Immersion liquids 1 new bottle of oil (Leica) 1 new bottle of glycerol (Leica) no

leica-dmrbe-med Changing of the control computer. Instigated by the DIT for 2013 (stop of win XP support). The camera uses now the Baumer drivers allowing to control it with Micromanager. no

leica-mz-16-fa-med Exchange of the control computer. Instigated by the DIT for 2013 (stop of win XP support) It's using the Baumer drivers now, allowing to control the leica camera with MicroManager. no

leica-tcs-sp5-med AOTF needs to be replaced. The ventilators broke down and some of the electronic components got damaged due to overheating no

leica-dm6000b-med Exchange of the Mercury lamp. 5600 hours. next lamp change at 6000 hours no

leica-tcs-sp5-med Argon laser and fibre (light guide) exchanged no

Why use it?

The Open Microscopy Environment aims to facilitate the handling of microscopy data as much as possible. For you this translates into the following features:

  • Organized and searchable database (find your data by experimenter, acquisition date, annotation…)
  • Easy access to your data: Connect with ImageJ, the OMERO Insight client or your web-browser
  • You can easily share your data with collaborators (single images, experiments, projects…)
  • No more image exporting. Instead of exporting the images from the microscopy software and loose all the meta-data, we import to OMERO.
  • Transparent metadata: Acquisition parameter like pixel-size, excitation wavelength, emission-wavelength etc. can be viewed at any time when browsing images
  • Easy vector graphics figure creation for article publication.
  • Facilitated and save transfer from the microscope on the server.
  • Facturing of the server space on the GB basis (as compared to 0.5TB by the DIT)
  • Fully integrated authentication: Users can login with their UniFr credentials


omero_web omero_general-infoomero_web_metadata
Web client (project  and dataset tree, dataset image
thumbnails annd image preview).
Interactive image viewer (virtual microscope).
vector graphics figure creation microscope image file medatata


Please login to read the user instructions